The question is find the value of "c" such that each pair of points is 5 units apart.
first pair: (5,2) (c,-3)
second pair: (0,c) (3,1)
Im not exactly sure what to do here. I haven't had a problem that has asked for an amount of units, not sure what units means. Would you possibly use the distance or midpoint formula to solve this, but again I don't know where the units part fits in...

Answer by stanbon(75887)

find the value of "c" such that each pair of points is 5 units apart.
first pair: (5,2) (c,-3)
distance = sqrt[(change in x)^2+(change in y)^2]
-----
5 = sqrt[(c-5)^2+(2--3)^2]
-----
5 = sqrt[c^2-10c+25 + 25]
-----
5 = sqrt[c^2-10c+50]
----
c^2-10c+50= 25
c^2-10c+25 = 0
(c-5)^2 = 0
c = 5
==============
second pair: (0,c) (3,1)
---
5 = sqrt[(3^2)+(c-1)^2]
---
5 = sqrt[9+c^2-2c+1]
25 = c^2-2c+10
c^2-2c-15 = 0
(c-5)(c+3) = 0
c = 5 of c= -3
=====================
Cheers,
Stan H.